8, Santos Unlimited (SU) was originally unlevered with 4000 shares outstanding. However, after a major financial restructure, SU now has $36000 of debt, with an annual interest expense of 6 percent. The restructuring has reduced the number of shares to 3000. A group of shareholders of SU are not convinced that this move towards adopting financial leverage is a good idea. Their main argument is that there is now some range of EBIT, however low, that will make the shareholders worse off than before. Help understand the situation better by computing the level of earnings before interest and tax (EBIT) that would make shareholders indifferent between being unlevered (i.e. not having any debt) and levered (i.e. having debt). Assume a 32 percent corporate tax rate.

Answer: $_____

Place your answer to the nearest dollar without a dollar sign or a comma (if applicable).
Answer #1

Let the level of earnings before interest and tax (EBIT) be EBITE that would make shareholders indifferent between being unlevered (i.e. not having any debt) and levered (i.e. having debt)

The EPS for both the firm will remain same at indifferent point

EPSunlevered = EPSlevered

PATunlevered/N1 = PATlevered/N2

(EBITE - I)(1-t)/N1 =(EBITE - I)(1-t)/N2

(EBITE - 0)/4000 =(EBITE - 36000*0.06)/3000

3000*EBITE = 4000*(EBITE - 2160)

1000*EBITE = 4000*2160

EBITE = $8640